# Low-resolution Prior Equilibrium Network for CT Reconstruction

### v1

@@ 1.2 Low-resolution Prior for CT reconstruction
The foundation of CT reconstruction methods is the Radon transform, which establishes a relationship between a 2D function $f(x,y)$ and a collection of line integrals [46]. The Radon transform of a parallel-beam imaging geometry for any arbitrary 2D object is illustrated in Figure 2. Let $\mathcal{L}(r,\varphi)$ represent the line in the Euclidean plane with an angle $\varphi$ and a signed distance $r$ from the origin
$\eqalign{\mathcal{L}(r,\varphi)&=\left\{(x,y)\in\mathbb{R}^{2}:x\cos\varphi+y\sin\varphi=r\right\}\cr&=\{(r\cos\varphi-\ell\sin\varphi,r\sin\varphi+\ell\cos\varphi):\ell\in\mathbb{R}\}.}$
Let $p_{\varphi}(r)$ denote the line of through $f(x,y)$ along the line $\mathcal{L}(r,\varphi)$
$p_{\varphi}(r)=\int_{\mathcal{L}(r,\varphi)}f(x,y)\mathrm{d}\ell=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}f(x,y)\delta(x\cos\varphi+y\sin\varphi-r)\mathrm{d}x\mathrm{d}y.$
Figure 2: Illustration of the CT imaging.
The Radon transform of $f$ is the complete collection of line integrals
$f\stackrel{{\scriptstyle Random}}{{\longleftrightarrow}}\{p_{\varphi}(r):\varphi\in[0,\pi],r\in(-\infty,\infty)\}.$ (5)
The function $p_{\varphi}(\cdot)$ is referred to as the projection of $f$ at angle $\varphi$. In its most idealized form, the problem of 2D image reconstruction involves recovering $f(x,y)$ from its projections {$p_{\varphi}(\cdot)$}. Here we use a 2D parallel beam geometry as an example to illustrate the discretization. As shown in Figure 3, let $N_{views}$ denote the number of views and $N_{bins}$ denote the finite number of radial samples. Then a total number of measured rays $\gamma_{i}=(\gamma_{1},\ldots,\gamma_{N_{views}\times N_{bins}})$ are recorded in the field of view (FOV), where $\Delta$ denotes the grid size used to approximate the continuous space. A smaller $\Delta$ corresponds to a higher sampling density in the spatial domain. The problem of reconstructing the object $f(x,y)$ from the measured sinogram $y$ is ill-posed due to the finite dimension of $\gamma$ and the infinite-dimensional nature of the unknown object $f(x,y)$. Suppose that the image $\bm{x}$ to be reconstructed from the sinogram $\gamma$ is of size $\bm{x}=(x_{1},\ldots,x_{N\times N})$. We can parameterize the object by utilizing a finite-series expansion [47]
$f(x,y)=\sum_{j=1}^{N\times N}x_{j}b_{j}(x,y),$ (6)
where $\{b_{j}(x,y)\}$ are basis functions that should be linearly independent. The square pixels are the most commonly used basis function defined by
$b_{j}(x,y)=rect\Big{(}\frac{x-\tilde{x}_{j}}{\Delta x}\Big{)}rect\Big{(}\frac{y-\tilde{y}_{j}}{\Delta y}\Big{)},$
@@
and ($\tilde{x}_{j}$,$\tilde{y}_{j}$) denotes the coordinates of the $j$-th pixel, ($\Delta x$, $\Delta y$) denotes the size of each pixel, typically square in shape. The length of the $i$th ray passing through the pixel $x_{j}$ is denoted by $a_{ij}=\int_{\mathcal{L}(r,\varphi)}b_{j}(x,y)\mathrm{d}\ell$, where all $a_{ij}$ constitute the system matrix $\bm{A}$ in (1).
Figure 3: Illustration of the virtual CT scanning system, where (a) a fine CT scanning system; (b) a coarse CT scanning system; (c) the limited-angle CT scanning system; (d) the sparse-view CT scanning system.
@@
In CT imaging, using a finer grid can improve image resolution, but it also leads to an increase in the number of unknowns, which exacerbates the ill-posedness of the inverse problem and increases sensitivity to disturbances [48]. We can achieve the goal of reducing the number of unknowns by increasing the size of pixels/voxels through modifying the basis functions (6). Actually, there have been studies on the impact of resolutions on CT reconstruction. In [49], a low-resolution full-view image was initially reconstructed, followed by a fine-resolution reconstruction of a specific region of interest (ROI). In particular, the image attributes of the pixels inside the ROI were set to zero, and a new projection sinogram was calculated using the forward model represented in matrix form by (1). Afterward, the newly calculated sinogram was subtracted from the original sinogram, resulting in a sinogram that contained projections exclusively corresponding to the ROI. Dabravolski et al. [50] proposed to obtain an initial reconstruction of a global coarse image, which was subsequently divided into regions comprising both fixed and non-fixed pixels. Specifically, the boundary pixels that shared at least one adjacent pixel with a different attenuation coefficient were assigned to the non-fixed pixels region. The non-fixed pixels region was comprised of boundary pixels that had at least one adjacent pixel with a different attenuation coefficient. Another method in [51] involved dividing the image domain into sub-regions with varying levels of discretization, which were characterized by either a coarse or fine pixel size. The projection data obtained from both resolutions were combined to reconstruct an image with different levels of discretization, effectively incorporating information from both measurements to enhance the reconstructed image quality. Gao et al. [52] proposed a low-resolution image prior image reconstruction model, which was realized in an end-to-end regime showing impressive ability in dealing with limited-angle reconstruction problems. He et al. [53] introduced a downsampling imaging geometric modeling approach for the data acquisition process, which integrated the geometric modeling knowledge of the CT imaging system with prior knowledge obtained from a data-driven training process to achieve precise CT image reconstruction. Malik et al. [54] suggested employing image fusion to generate a composite image that combines the high spatial resolution of a partially ambiguous image obtained from incomplete data with the more dependable quantitative information of a coarser image reconstructed from the same data in an over-complete problem. The aforementioned methods demonstrate that the low-resolution image can be used as an effective priori for CT reconstruction.
1.3 Our Contribution
@@
In this work, we introduce the low-resolution image as the prior and propose a new energy minimization problem for CT reconstruction. The low-resolution image prior can facilitate the reconstruction quality, especially for incomplete data such as sparse-view and limited-angle reconstruction. Subsequently, we implement the algorithm unrolling to solve our low-resolution prior image model, where the gradient descent is employed for minimizing the lower-level problem with each Jacobi block being approximated by convolutional neural networks. To balance feature extraction capability and model size, we establish the low-resolution prior equilibrium (LRPE) network, where the weight-sharing strategy is used among all iterations. More importantly, our approach has provable convergence guarantees by satisfying certain conditions to guarantee the iterates converge to a fixed-point. Through extensive numerical experiments on both sparse-view and limited-angle CT reconstruction problems, our LRPE demonstrates clear advantages over the state-of-the-art learning methods. To sum up, the major contributions of this paper are as follows
@@
  • By leveraging the characteristics of the physical process of CT imaging, we use the low-resolution image as prior and establish a novel energy minimization model.
@@ 2 Our Low-resolution Prior Reconstruction Model
Let $\bm{A_{l}}$ be the system matrix of size $(N_{views}\times N_{bins})\times(n\times n)$, and $\bm{x_{l}}$ be the low-resolution image of size $n\times n$. Then the corresponding inverse problem can be formulated as follows
$\bm{A_{l}}\bm{x_{l}}=\bm{b},$ (7)
@@
where $n<N$ denotes the length and width of the low-resolution image. Without considering the discretization error, the same projection data $\bm{b}$ is used for the low-resolution image. The visual comparison between the fine and coarse CT scanning system is provided in the first row of Figure 3. As can be observed, the red color dots denote the permissible positions for the emitter, and the blue arc represents the corresponding receiving range of the receiver. It is obviously shown that the coarse scanning system has the same field of view (FOV) but with a larger step size in discretizing the FOV. Thus, the low-resolution image can be obtained from the high-resolution image as follows
$\bm{x_{l}}=\bm{D}\bm{x},$ (8)
where $D$ denotes the down-sampling opterator of size $(n\times n)\times(N\times N)$. For the given data $\bm{b}\in\mathbb{R}^{N_{views}\times N_{bins}}$, we consider the following lower-level problem
${\bm{x}}\in\mathop{\arg\min}\limits_{\bm{x}}\mathcal{S}(\bm{Ax},\bm{b})+\mathcal{R}(\bm{x},\bm{D^{\top}x_{l}}),$ (9)
@@
where $\bm{D^{\top}}$ represents the transpose of the matrix $\bm{D}$. Supposing both $\mathcal{S}$ and $\mathcal{R}$ are differentiable, we can use the gradient descent to solve the minimization problem (9). That is, we start with an initial estimate $\bm{x}^{(0)}$ such as $\bm{x}^{(0)}=0$ and choose a step size $\eta>0$, such that for iteration $k=1,2,3,\ldots$, we can obtain
$\bm{x}^{(k+1)}=\bm{x}^{(k)}-\eta\bm{A^{\top}}\nabla_{{x}}{\mathcal{S}}(\bm{Ax}^{(k)},\bm{b})-\eta\nabla_{{x}}\mathcal{R}\left(\bm{x}^{(k)},\bm{D^{\top}x_{l}}\right),$
where $\nabla_{x}\mathcal{S}$ and $\nabla_{x}\mathcal{R}$ represent the gradients of the data term and the regularization term, respectively. The idea behind unrolling strategy is to select a predetermined number of iterations, denoted as $N$, and designate $\bm{x}^{(N-1)}$ as our estimated value ${\bm{x}}$. We employ neural networks $S_{\theta}(\bm{x})$ and $R_{\theta}(\bm{x})$ to represent $\nabla_{x}\mathcal{S}$ and $\nabla_{x}\mathcal{R}$, respectively, the weights of which are denoted as $\bm{\theta}$ to be learned from the training data. More specifically, we define the unrolled gradient descent as ${\bm{x}}^{(N-1)}(\bm{b};\bm{\theta}):=\bm{x}^{(N-1)}$ with $\bm{x}^{(0)}=0$. For $k=0,\ldots,N-1$, we obtain the recursive update as follows
$\bm{x}^{(k+1)}=\bm{x}^{(k)}-\eta\bm{A^{\top}}S_{\theta}(\bm{Ax}^{(k)},\bm{b})-\eta R_{\theta}(\bm{x}^{(k)},\bm{D^{\top}x_{l}}).$ (10)
In favor of the deep equilibrium model in improving feature abstraction capabilities by unrolling the network to a sufficient depth, we use the weight-sharing strategy to design an iterative map $f_{\theta}(\cdot;\bm{b})$ that leads to a fixed-point $\bm{x}^{(\infty)}$ satisfying
$\bm{x}^{(\infty)}=f_{\theta}\left(\bm{x}^{(\infty)};\bm{b}\right).$ (11)
Connecting the unrolled gradient descent iterations (10) with the deep equilibrium architecture (4), we introduce the following formulation
$f_{\theta}(\bm{x};\bm{b})=\bm{x}-\eta\bm{A}^{\top}S_{\theta}(\bm{Ax},\bm{b})-\eta R_{\theta}(\bm{x},\bm{D^{\top}x_{l}}).$ (12)
@@
The objective in the upper-level problem is simply the true risk for the squared loss
$\ell(\bm{{{x}}^{*}},\bm{x})=\frac{1}{2}\|\bm{{{x}}^{*}}-\bm{x}\|_{2}^{2},$
@@
where $\bm{x}$ denotes the reconstructed image, and $\bm{x}^{*}$ denotes the ground truth. Indeed, the use of the mean squared error (MSE) loss is not necessary, as any differentiable loss function can be used.
@@ 3 Algorithm Implementation
Our end-to-end Low-Resolution Prior Equilibrium (LRPE) network is described in Algorithm 1, which is implemented in Python using the Operator Discretization Library (ODL), the Adler package, the ASTRA Toolbox, and TensorFlow 1.8.0. The tensorFlow is a toolkit designed for tackling complex mathematical problems, for which the calculations are represented as graphs, mathematical operations are nodes, and multidimensional data arrays are communicated as edges of the graphs. The ASTRA Toolbox is a MATLAB and Python toolbox that provides high-performance GPU primitives for 2D and 3D tomography. And ODL is a Python library focused on fast prototyping for inverse problems. Adler is a toolkit that facilitates the efficient implementation of neural network architectures.
@@
1:Step 0. Initialize $\bm{x}^{0}$, $\bm{b}$, $\bm{x_{l}}$, step size ${\eta}$;
2:Step 1. For $k=0,1,2,\cdots,N-1$ do
$\left\{\begin{array}[]{l}R_{\theta}^{k+1}\longleftarrow\Gamma_{\theta}(\bm{x}^{k},\bm{D^{\top}x_{l}});\\ S_{\theta}^{k+1}\longleftarrow\Lambda_{\theta}(\bm{Ax}^{k},\bm{b});\\ \bm{x}^{k+1}=\bm{x}^{k}-\eta(\bm{A}^{\top}S_{\theta}^{k+1}+R_{\theta}^{k+1});\\ \end{array}\right.$
3:Step 2. Return $\bm{x}^{N-1}$.

### v2

@@ 1.2 Low-resolution Prior for CT reconstruction
For severely degraded CT reconstruction problems such as sparse-view and limited-angle, it is necessary to combine effective prior information to improve the quality of the reconstructed images. Different kinds of prior knowledge have been investigated for incomplete data reconstruction problems. The non-local similarity has been shown highly beneficial for enhancing fine details in reconstructed CT images [32, 35, 21, 45]. Both low-rank prior and dictionary learning have been used to leverage the sparsity of the image in a transform domain [41, 11, 53]. Deep learning methods by integrating geometric priors have also been studied in [58, 39, 64, 71]. Besides, personalized information as a prior has also been demonstrated to significantly enhance the reconstruction results [20, 24].
In CT imaging, the problem of reconstructing the object $\bm{u}$ from the measured sinogram $\bm{b}$ is ill-posed due to the finite dimension of measured rays and the infinite-dimensional nature of the unknown object $\bm{u}$. We can parameterize the object by utilizing a finite-series expansion as follows
$\bm{u}(x,y)=\sum_{j=1}^{N^{2}}u_{j}e_{j}(x,y),$ (6)
where $\{e_{j}(x,y)\}$ are basis functions that should be linearly independent. The square pixels are the most commonly used basis function defined by
$e_{j}(x,y)=rect\Big{(}\frac{x-\tilde{x}_{j}}{\Delta x}\Big{)}rect\Big{(}\frac{y-\tilde{y}_{j}}{\Delta y}\Big{)},$
@@
and ($\tilde{x}_{j}$,$\tilde{y}_{j}$) denotes the coordinates of the $j$-th pixel, ($\Delta x$, $\Delta y$) denotes the size of each pixel. The length of the $i$th ray passing through the pixel $u_{j}$ is denoted by $a_{ij}=\int e_{j}(x,y)\mathrm{d}\ell$, where all $a_{ij}$ constitute the system matrix $\bm{A}$ in (1). It is well-known that using a finer grid can improve image resolution, but it also leads to an increase in the number of unknowns, which exacerbates the ill-posedness of the inverse problem and increases sensitivity to disturbances [27]. We can achieve the goal of reducing the number of unknowns by increasing the size of pixels/voxels through modifying the basis functions (6). Actually, there have been studies on the impact of resolutions on CT reconstruction. In [62], a low-resolution full-view image was initially reconstructed, followed by a fine-resolution reconstruction of a specific region of interest (ROI). In particular, the image attributes of the pixels inside the ROI were set to zero, and a new projection sinogram was calculated using the forward model represented in matrix form by (1). Afterward, the newly calculated sinogram was subtracted from the original sinogram, resulting in a sinogram that contained projections exclusively corresponding to the ROI. Dabravolski et al. [14] proposed to obtain an initial reconstruction of a global coarse image, which was subsequently divided into regions comprising both fixed and non-fixed pixels. Specifically, the boundary pixels that shared at least one adjacent pixel with a different attenuation coefficient were assigned to the non-fixed pixels region. The non-fixed pixels region was comprised of boundary pixels that had at least one adjacent pixel with a different attenuation coefficient. Another method in [9] involved dividing the image domain into sub-regions with varying levels of discretization, which were characterized by either a coarse or fine pixel size. The projection data obtained from both resolutions were combined to reconstruct an image with different levels of discretization, effectively incorporating information from both measurements to enhance the reconstructed image quality. Gao et al. [17] proposed a low-resolution image prior image reconstruction model, which was realized in an end-to-end regime showing impressive ability in dealing with limited-angle reconstruction problems. He et al. [29] introduced a downsampling imaging geometric modeling approach for the data acquisition process, which integrated the geometric modeling knowledge of the CT imaging system with prior knowledge obtained from a data-driven training process to achieve precise CT image reconstruction. Malik et al. [42] suggested employing image fusion to generate a composite image that combines the high spatial resolution of a partially ambiguous image obtained from incomplete data with the more dependable quantitative information of a coarser image reconstructed from the same data in an over-complete problem. The aforementioned methods demonstrate that the low-resolution image can be used as an effective prior for ill-posed CT reconstruction problems.
1.3 Our contributions
@@
In this work, we introduce the low-resolution image as the prior and propose a novel image reconstruction model for sparse-view and limited-angle CT reconstruction tasks. More specifically, the low-resolution image prior can introduce effective regularity into the reconstruction model to guarantee the qualities for incomplete data problems. Subsequently, we implement the algorithm unrolling to solve our low-resolution prior image model, where the gradient descent is employed for minimizing the lower-level problem with each Jacobi block being approximated by convolutional neural networks. To balance feature extraction capability and model size, we establish the low-resolution prior equilibrium (LRPE) network, where the weight-sharing strategy is used among all stages. More importantly, our approach has provable convergence guarantees by satisfying certain conditions to guarantee the iterative scheme converge to a fixed-point. Through extensive numerical experiments on both sparse-view and limited-angle CT reconstruction problems, our LRPE model demonstrates clear advantages over the state-of-the-art learning methods. To sum up, the major contributions of this paper are as follows
@@
  • By leveraging the characteristics of the physical process of CT imaging, we use the low-resolution image as prior and establish an effective regularization model for incomplete data CT reconstruction.
@@
Figure 4: Illustration of the virtual CT scanning system, where (a) a fine CT scanning system; (b) a coarse CT scanning system.
@@ 2 Our Low-resolution Prior Reconstruction Model
Let $\bm{A_{l}}:\mathbb{R}^{n^{2}}\rightarrow\mathbb{R}^{N_{views}\times N_{bins}}$ be the system matrix for the low-resolution image $\bm{u_{l}}$. Thus, the low-resolution inverse problem can be formulated as follows
$\bm{A_{l}}\bm{u_{l}}=\bm{b},$ (7)
@@
where $n$ denotes the length and width of the low-resolution image. Without considering the discretization error, the same projection data $\bm{b}$ is used for the low-resolution image. The illustration of the fine and coarse CT scanning system is provided in Figure 4. As can be observed, the red color dots denote the permissible positions for the emitter, and the blue arc represents the corresponding receiving range of the receiver. It is obviously shown that the coarse scanning system has the same field of view (FOV) but with a larger step size in spatial discretization. That means the low-resolution image can be obtained from the high-resolution image as follows
$\bm{u}_{l}=\bm{D}\bm{u},$ (8)
where $\bm{D}:\mathbb{R}^{N^{2}}\rightarrow\mathbb{R}^{n^{2}}$ is the down-sampling operator. In this work, we concern with the following low-resolution image regularized CT reconstruction model
${\bm{u}}\in\mathop{\arg\min}\limits_{\bm{u}}\mathcal{S}(\bm{Au},\bm{b})+\mathcal{R}(\bm{u},\bm{D^{\top}u}_{l}),$ (9)
@@
where $\bm{D^{\top}}$ represents the transpose of the matrix $\bm{D}$. Supposing both $\mathcal{S}$ and $\mathcal{R}$ are differentiable, we can use the gradient descent to solve the minimization problem (9). That is, we start with an initial estimate $\bm{u}^{(0)}$ such as $\bm{u}^{(0)}=0$ and a step size $\eta>0$, such that for stage $k=0,1,\ldots,K-1$, we estimate $\bm{u}^{(k+1)}$ from
$\bm{u}^{(k+1)}=\bm{u}^{(k)}-\eta\Big{(}\bm{A^{\top}}\nabla{\mathcal{S}}\left(\bm{Au}^{(k)},\bm{b}\right)+\nabla\mathcal{R}\left(\bm{u}^{(k)},\bm{D^{\top}u}_{l}\right)\Big{)}.$
To leverage the rich information contained in the data, we adopt the deep unrolling strategy to solve the above gradient descent problem. More specifically, we utilize two convolutional neural networks $R_{\theta}(\cdot,\cdot):\mathbb{R}^{N^{2}}\times\mathbb{R}^{N^{2}}\rightarrow\mathbb{R}^{N^{2}}$ and $S_{\theta}(\cdot,\cdot):\mathbb{R}^{N_{views}\times N_{bins}}\times\mathbb{R}^{N_{views}\times N_{bins}}\rightarrow\mathbb{R}^{N_{views}\times N_{bins}}$ to learn the gradient operator $\nabla\mathcal{R}\left(\bm{u}^{(k)},\bm{D^{\top}u_{l}}\right)$ and $\nabla\mathcal{S}\left(\bm{Au}^{(k)},\bm{b}\right)$, respectively. For $k=0,\ldots,K-1$, we then have the recursive update as follows
$\bm{u}^{(k+1)}=\bm{u}^{(k)}-\eta(\bm{A^{\top}}s^{k}+t^{k}),$ (10)
where $s^{k}:=S_{\theta}(\bm{Au}^{(k)},\bm{b})$ and $t^{k}:=R_{\theta}(\bm{u}^{(k)},\bm{D^{\top}u_{l}})$ are the estimations of the two convolutional neural network models, respectively. Moreover, we use the weight-sharing scheme to improve the feature abstraction capabilities by unrolling the network to a sufficient depth. Therefore, the final update scheme is achieved by connecting the unrolled gradient descent stages (10) with the deep equilibrium architecture (5), as follows
$f_{\theta}(\bm{u}^{k+1};\bm{b}):=\bm{u}^{k}-\eta\Big{(}\bm{A}^{\top}S_{\theta}(\bm{Au^{k}},\bm{b})+R_{\theta}(\bm{u}^{k},\bm{D^{\top}u}_{l})\Big{)}.$ (11)
@@
The objective function in the upper-level problem is simply the true risk for the squared loss
$\ell(\bm{u},\bm{u}^{*})=\frac{1}{m}\sum_{i=1}^{m}\|{{{\bm{u}_{i}}}^{*}}-{\bm{u}_{i}}^{K}\|_{2}^{2},$
@@
where ${\bm{u}_{i}}^{*}$ is the $i$th ground truth, and ${\bm{u}}^{K}_{i}$ denotes the reconstructed image. Indeed, the use of the mean squared error (MSE) loss is not necessary, as any differentiable loss function can be used.
@@ 3 Algorithm Implementation
The low-resolution prior equilibrium (LRPE) model can be outlined as Algorithm 1. In both training and inference stage, the low-resolution prior image $\bm{u}_{l}$ is obtained by established reconstruction methods, where we used the Learned Primal-Dual network [1] to estimate the low-resolution image in our work.
@@
1:Step 0. Initialize $\bm{u}^{0}$, $\bm{b}$, $\bm{u_{l}}$, step size ${\eta}$;
2:Step 1. For $k=0,1,2,\cdots,K-1$ do
$\left\{\begin{array}[]{l}t^{k}\longleftarrow S_{\theta}(\bm{u}^{k},\bm{D^{\top}u_{l}});\\ s^{k}\longleftarrow R_{\theta}(\bm{Au}^{k},\bm{b});\\ \bm{u}^{k+1}=\bm{u}^{k}-\eta(\bm{A}^{\top}s^{k}+t^{k});\\ \end{array}\right.$
3:Step 2. Return $\bm{u}^{K}$.
In implementation, our algorithm was realized using the Operator Discretization Library (ODL), the Adler package, the ASTRA Toolbox, and TensorFlow 1.8.0. In particular, the TensorFlow is a toolkit designed for tackling complex mathematical problems, for which the calculations are represented as graphs, mathematical operations are nodes, and multidimensional data arrays are communicated as edges of the graphs. The ASTRA Toolbox is a MATLAB and Python toolbox that provides high-performance GPU primitives for 2D and 3D tomography. And ODL is a Python library focused on fast prototyping for inverse problems. Adler is a toolkit that facilitates the efficient implementation of neural network architectures.